# Prediction of Protein Three-dimensional Structures via a Hardware-Executable Quantum Computing Framework

### v1

@@ A Quantum Engineering Framework for Structure Prediction
Here we select a representative experimental example to further illustrate the results. As shown in Figure 3B-E, we chose a fragment from the binding pocket region of protein 3b26 for demonstration. In our experiments, this fragment consists of 11 amino acids—one of the longer test cases—which facilitates clear visualization. In the visualized outcome, the structure predicted by the quantum method clearly and accurately captures the overall fold and key structural features of the selected segment, showing a high degree of similarity to the experimentally determined structure in the dataset. In contrast, the AF3 prediction shown in Figure E deviates substantially from the experimental result: AF3 models this region as an $\alpha$-helical secondary structure, which differs markedly from the true conformation and loses essential structural characteristics. In our experiments, many AF3 predictions tended to default to helical structures for unknown regions. We hypothesize that this bias arises because helices have high recognizability and pronounced patterns, which create an information “pit” in pattern-recognition models; under limited system information capacity, the deep learning model gravitates toward predictions with the most distinctive features, thus sacrificing accuracy. In contrast, quantum methods build from first principles and do not fall into pattern-recognition traps. Even when system information capacity is severely constrained, they can still produce results with reliable fidelity.
@@
Figure 3: Structural testing and visual representation (A) The docking workflow. The predicted structures from both methods undergo hydrogen addition and charge neutralization, followed by docking tests with the same receptor molecule from the dataset. To ensure fairness in evaluation, both prediction results are processed using an identical postprocessing pipeline. (B) Native structure of 3b26. (C) Docking pocket view. (D) Quantum overlay. The quantum method essentially captures the overall trajectory of the regional fragment and the backbone amino acid coordinates, showing high consistency with the reference (scanned) structure. (E) AF3 overlay. The prediction from AF3 appears to have fallen into a locally optimal state, exhibiting a tendency to converge toward structurally prominent features when the available information capacity (i.e., amino acid chain length) is limited. (F) RMSD comparison. Comparison of the structural differences in RMSD between the predictions of the two methods and the reference structures from the dataset. A smaller RMSD value indicates that the predicted structure is closer to the actual structure, reflecting higher prediction accuracy. (G) Affinity comparison. (H) RMSD upper bound comparison. (I) RMSD lower bound comparison.
@@ Discussion
As protein sequences lengthen and the number of interactions increases, the corresponding Hamiltonian grows more complex, and the depth of the quantum circuit also increases. Consequently, noise accumulation becomes more pronounced, which can slow VQE convergence, trap the system in local minima, or shift the estimated energy away from its true value, all of which undermine the accuracy of the protein backbone conformation. Meanwhile, if the number of measurements in the final sampling stage is insufficient, statistical (shot) noise remains large; if measurements are too frequent, hardware is required to maintain the same circuit state for a longer duration, possibly amplifying noise further [35]. Some studies have demonstrated that using coarse-grained or lower-dimensional energy functions to represent protein structures can reduce the reliance on fully atomistic simulations and significantly decrease the sensitivity of quantum systems to noise [36, 37, 38]. Meanwhile, recent research provides further evidence that non-fault-tolerant quantum devices still possess sufficient performance to address utility tasks [39]. In our work, we employed a simplified vector-based protein backbone model, which has been shown to exhibit high robustness [40, 41], reducing the dimensionality of observables and the overall state space. Although this coarse-grained model cannot perfectly capture atomic-scale interactions, it is sufficient to represent key geometric features and energy trends. Our five model workflow also significantly reduces the impact of noise. In the first stage (estimation), the system avoids large-scale quantum-state measurements and focuses on rapidly estimating total energy, thereby reducing the accumulation of quantum gate and readout errors. In the second stage (sampling), after the circuit parameters have approximately converged, we perform extensive measurements in a single run to fully extract the final quantum-state distribution, applying readout error mitigation techniques to correct for systematic biases in the measurement process. While utility-level quantum devices still exhibit disturbances caused by noise, these disturbances do not have a practical impact on the results or significantly affect the overall performance of the system in the problem scenarios we addressed (Figures 2 and 4).
@@
To mitigate noise in quantum processors, previous studies have proposed a range of techniques, such as adding error detection or error correction codes (ECC) to quantum circuits [42, 43, 44], or deploying denoisers at the pulse or measurement level [45, 46]. While these measures effectively reduce error rates, they also require additional physical qubits and quantum gate operations, increasing hardware demands and runtime [47]. Additionally, zero-noise extrapolation (ZNE) artificially amplifies circuit noise by inserting pairs of quantum gate operations, measures energy values under different noise levels, and extrapolates results toward a ”zero-noise” limit  [48]. Pulse-level optimizations and dynamic error correction may further suppress decoherence [49]. As hardware evolves and quantum error-correcting codes (QEC) mature, gate fidelities and the number of available qubits are expected to increase, providing broader possibilities for larger-scale and higher-accuracy protein structure prediction [50, 51, 52, 53]. Looking ahead, advancements in quantum hardware, including increased qubit numbers and improved error-correction capabilities, will enable quantum computing to address more complex biological tasks. Potential application areas include long-chain protein structure prediction, modeling multi-molecular interactions, and drug molecule design. Moreover, hybrid approaches combining classical optimizers with deep learning algorithms may further enhance the efficiency and applicability of quantum computing, driving revolutionary advancements in computational biology.
@@ Methods
The data used in this study were obtained from the PDBbind database [11, 12, 13], which contains high-quality protein-ligand complex structures and protonation charge data. Protein structures with well-defined active sites were selected to ensure consistent structural information and charge states, thereby standardizing the inputs for the prediction models. Because our approach aims to provide a general solution for protein-ligand docking problems, rather than focusing on any specific drug discovery target, we randomly selected two sets of complexes from the PDBbind database for our experiments. There was no particular rationale for choosing these two sets beyond their availability and completeness; they serve as representative examples to demonstrate the universality of our method. The PDBbind database is widely accepted as a standard resource for evaluating docking and binding-affinity prediction methods and has been used in numerous previous studies [54, 55, 56, 57].
@@
We adopted a tetrahedral lattice modeling approach inspired by previous studies [58] to construct the geometric backbone of proteins from amino acid sequences. Each amino acid node has four possible extension directions, represented by a set of directional vectors[59]. The distance between adjacent residues is fixed to unit length, and the bond angle is set to approximately 109.4°, consistent with fundamental protein geometry. This coarse-grained representation preserves the essential structural features of proteins while significantly reducing the number of required qubits. In our approach, the connection vectors between amino acid nodes, inter-residue interactions, and structural constraints are all mapped to a set of quantum operators. During computation, the number of qubits required by the quantum processor is determined by the dimensionality of these operators derived from the modeling process. In quantum systems, each qubit encodes a $2$-dimensional Hilbert space, so $n$ qubits can represent a space of dimension $2^{n}$. Once the system Hamiltonian is defined, the processor must allocate sufficient qubits to span the solution space. The number of qubits increases with problem complexity, which is jointly determined by the variety and length of the input amino acid sequence. Amino acid types influence the scale of interaction and constraint operators, while sequence length directly determines the dimensionality of the solution space. Table 1 summarizes the number of qubits and circuit depths used in each case. The core modeling logic exploits the superposition property of quantum systems. For example, if we encode an amino acid node $A$ using 4 qubits, each measurement outcome (e.g., 0001, 0010, 0100, or 1000) activates one of four extension vectors corresponding to directions in the tetrahedral lattice[60].. At the start, all possible connection paths for the sequence are encoded in superposition, where each amino acid node has four equally probable directions. Under the influence of the Hamiltonian, the variational optimization gradually alters the probability distribution of each path. During measurement, the quantum state collapses to the most probable configuration, represented by a binary string in which only one direction is activated per node. This outcome corresponds to the most likely amino acid connection determined by energy minimization. Since the Hamiltonian encodes constraints as energy penalties, invalid configurations lead to higher energy terms and are suppressed during optimization. As a result, the energy landscape of the protein is reflected in the energy domain of the quantum system, allowing quantum computation to efficiently search the conformational space via its inherent superposition capability. The quantum state evolves toward the conformation that minimizes the protein’s total energy.
Hamiltonian Design
The total Hamiltonian for protein structure prediction, $H_{\text{t}}$, consists of four main components:
$H_{\text{t}}=H_{\text{c}}+H_{\text{g}}+H_{\text{d}}+H_{\text{i}}$
$H_{\text{c}}$: Penalizes configurations that violate the natural chirality of amino acids, ensuring stereochemical correctness. $H_{\text{g}}$: Imposes geometric constraints to maintain backbone integrity. $H_{\text{d}}$: Penalizes unreasonable distance interactions, preventing spatial collisions. $H_{\text{i}}$: Encodes interaction energies between amino acids to identify the lowest-energy structure[61, 62, 63].
By minimizing $H_{\text{t}}$, the quantum system searches for the optimal protein conformation that satisfies both geometric and physical constraints. We utilized the Variational Quantum Eigensolver (VQE) to identify the ground state of the system and efficiently encoded the Hamiltonian using the SparsePauliOp class in Qiskit. The EfficientSU2 Ansatz was chosen for its balance between expressiveness and robustness against device noise, while the COBYLA algorithm served as the classical optimizer, suitable for utility-level environments.
We further introduced weighting factors $\lambda_{c},\lambda_{g},\lambda_{d},$ and $\lambda_{i}$ (all set to 1 in our current implementation) to adjust the relative strength of each term:
$H_{\text{t}}=\lambda_{c}H_{\text{c}}+\lambda_{g}H_{\text{g}}+\lambda_{d}H_{\text{d}}+\lambda_{i}H_{\text{i}}.$
These weights can be tuned to emphasize particular aspects of the protein structure, such as tighter chirality constraints or stronger inter-residue interactions. In our experiments, we found that setting all $\lambda$’s to the same value provided a reasonable balance for diverse protein sequences. However, specific cases may benefit from adjusting these coefficients, especially when dealing with atypical residues or post-translational modifications.
@@
In realistic scenarios, the qubit requirement grows significantly with the amino acid sequence length. For very short sequences, fewer than 10 qubits may suffice. However, in our experiments, when dealing with 12 amino acids, at least 77 qubits are needed to encode all relevant operators. Moreover, to satisfy certain coupling conditions and allow for additional SWAP operations that reduce overall circuit depth, we reserve about 5 extra qubits (totaling 82) as a buffer. This strategy effectively lowers the entangling-gate overhead (see, e.g., [64, 65] for related approaches). In many recent studies, the EfficientSU2 ansatz has emerged as a widely used and robust parameterized circuit, recognized for its relatively broad applicability and moderate resource requirements [66, 67, 68]. To parameterize our circuits, we employ the EfficientSU2 ansatz from Qiskit. This ansatz typically consists of repeated blocks of single-qubit rotation gates $R_{y}$ and $R_{z}$, interspersed with CX gates arranged in a circular or linear topology to introduce entanglement. By adjusting the number of repetition layers (reps), we can flexibly control the circuit’s expressive power. In a single repetition layer, each qubit undergoes one $R_{y}R_{z}$ rotation, followed by an entangling step that couples adjacent qubits. The theoretical circuit depth of Qiskit’s EfficientSU2 parameterized quantum circuit grows linearly with the number of qubits $N$. For our case with $N=82$ qubits: Total theoretical depth is $333$. The actual compiled depth may be different through gate optimization and hardware topology adaptation. An important advantage of the quantum method is its extremely fast computational speed. Due to the superposition state of the quantum system, during each optimization, the system does not focus on a specific structure for validation like traditional structure-based prediction methods. Instead, the superposition state allows all possible conformations of the current protein structure to exist simultaneously at the physical level, collapsing into a definite ground state as the optimal result through parameter optimization. In our tests, for a sequence with a length of 5 amino acids, the total recorded execution time after 150 iterations was 4079.55 seconds. According to logs from IBM’s superconducting quantum computer, the average time per optimization iteration and state measurement, from task submission to result acquisition, was approximately 10 seconds, with about 73.53% of the time spent on the quantum end. In our setup, the number of shots for optimization tasks was 200, and the number of shots for measurement was 100,000, indicating that the execution efficiency of the quantum computing method is extremely high. The quantum processor itself completes a single task almost instantaneously, with most of the time spent on state preparation and the machine’s self-calibration. Additionally, part of the time was limited by the performance of the classical computing backend. Under computational models of equivalent accuracy, the quantum method is several orders of magnitude faster than all current methods based on classical computing.

### v2

@@ A Quantum Engineering Framework for Structure Prediction
Here we select a representative experimental example to further illustrate the results. As shown in Figure 3B-E, we chose a fragment from the binding pocket region of protein 3b26 for demonstration. In our experiments, this fragment consists of 11 amino acids—one of the longer test cases—which facilitates clear visualization. In the visualized outcome, the structure predicted by the quantum method clearly and accurately captures the overall fold and key structural features of the selected segment, showing a high degree of similarity to the experimentally determined structure in the dataset. In contrast, the AF3 prediction shown in Figure E deviates substantially from the experimental result: AF3 models this region as an $\alpha$-helical secondary structure, which differs markedly from the true conformation and loses essential structural characteristics. The classical model, by contrast, produces a highly distorted conformation that fails to preserve backbone geometry or side-chain orientation, consistent with its higher RMSD and weaker docking affinity (Figure 3F–G). In our experiments, many AF3 predictions tended to default to helical structures for unknown regions. We hypothesize that this bias arises because helices have high recognizability and pronounced patterns, which create an information “pit” in pattern-recognition models; under limited system information capacity, the deep learning model gravitates toward predictions with the most distinctive features, thus sacrificing accuracy. In contrast, quantum methods build from first principles and do not fall into pattern-recognition traps. Even when system information capacity is severely constrained, they can still produce results with reliable fidelity.
@@
Figure 3: Structural testing and visual representation (A) Docking workflow. Predicted structures from both methods are hydrogenated, charge-neutralized, and docked against the same receptor using an identical postprocessing pipeline. (B) Native structure of 3b26. (C) Docking pocket view. (D) Quantum overlay. The quantum prediction accurately follows the fragment trajectory and backbone conformation, showing high consistency with the reference structure. (E) AF3 overlay. The AF3 prediction tends to converge to a local minimum under limited sequence information. (F) RMSD comparison. Structural deviation between predicted and reference structures; smaller RMSD indicates higher accuracy. (G) Affinity comparison. (H) RMSD upper bound. (I) RMSD lower bound.
@@ Discussion
As protein sequences lengthen and the number of interactions increases, the corresponding Hamiltonian grows more complex, and the depth of the quantum circuit also increases. Consequently, noise accumulation becomes more pronounced, which can slow VQE convergence, trap the system in local minima, or shift the estimated energy away from its true value, all of which undermine the accuracy of the protein backbone conformation. Meanwhile, if the number of measurements in the final sampling stage is insufficient, statistical (shot) noise remains large; if measurements are too frequent, hardware is required to maintain the same circuit state for a longer duration, possibly amplifying noise further [50]. Some studies have demonstrated that using coarse-grained or lower-dimensional energy functions to represent protein structures can reduce the reliance on fully atomistic simulations and significantly decrease the sensitivity of quantum systems to noise [80]. Meanwhile, recent research provides further evidence that non-fault-tolerant quantum devices still possess sufficient performance to address utility tasks [33, 46]. In our work, we employed a simplified vector-based protein backbone model, which has been shown to exhibit high robustness [14, 84], reducing the dimensionality of observables and the overall state space. Although this coarse-grained model cannot perfectly capture atomic-scale interactions, it is sufficient to represent key geometric features and energy trends. Our five model workflow also significantly reduces the impact of noise. In the first stage (estimation), the system avoids large-scale quantum-state measurements and focuses on rapidly estimating total energy, thereby reducing the accumulation of quantum gate and readout errors. In the second stage (sampling), after the circuit parameters have approximately converged, we perform extensive measurements in a single run to fully extract the final quantum-state distribution, applying readout error mitigation techniques to correct for systematic biases in the measurement process. While utility-level quantum devices still exhibit disturbances caused by noise, these disturbances do not have a practical impact on the results or significantly affect the overall performance of the system in the problem scenarios we addressed (Figures 2 and 4).
@@
To mitigate noise in quantum processors, previous studies have proposed a range of techniques, such as adding error detection or error correction codes (ECC) to quantum circuits [37], or deploying denoisers at the pulse or measurement level [19]. While these measures effectively reduce error rates, they also require additional physical qubits and quantum gate operations, increasing hardware demands and runtime [70]. Additionally, zero-noise extrapolation (ZNE) artificially amplifies circuit noise by inserting pairs of quantum gate operations, measures energy values under different noise levels, and extrapolates results toward a ”zero-noise” limit  [23]. Pulse-level optimizations and dynamic error correction may further suppress decoherence [28]. As hardware evolves and quantum error-correcting codes (QEC) mature, gate fidelities and the number of available qubits are expected to increase, providing broader possibilities for larger-scale and higher-accuracy protein structure prediction [83, 17, 25]. Looking ahead, advancements in quantum hardware, including increased qubit numbers and improved error-correction capabilities, will enable quantum computing to address more complex biological tasks. Potential application areas include long-chain protein structure prediction, modeling multi-molecular interactions, and drug molecule design. Moreover, hybrid approaches combining classical optimizers with deep learning algorithms may further enhance the efficiency and applicability of quantum computing, driving revolutionary advancements in computational biology.
@@ Methods
The data used in this study were obtained from the PDBbind database [73, 74, 44], which contains high-quality protein-ligand complex structures and protonation charge data. Protein structures with well-defined active sites were selected to ensure consistent structural information and charge states, thereby standardizing the inputs for the prediction models. Because our approach aims to provide a general solution for protein-ligand docking problems, rather than focusing on any specific drug discovery target, we randomly selected two sets of complexes from the PDBbind database for our experiments. There was no particular rationale for choosing these two sets beyond their availability and completeness; they serve as representative examples to demonstrate the universality of our method. The PDBbind database is widely accepted as a standard resource for evaluating docking and binding-affinity prediction methods and has been used in numerous previous studies [62, 58, 68].
@@
A central principle of our encoding strategy is the control of structural granularity when projecting protein conformations into a quantum representation. Earlier lattice-based approaches [55, 18, 85] typically employed link- or vector-based encodings that preserve both backbone and side-chain variability [82]. While such fine-grained models capture detailed features, they also lead to steep, rugged energy landscapes that amplify hardware noise and increase the qubit requirements of the quantum system. In this work, we introduce a coarser-grained, single-chain vector representation that omits side-chain degrees of freedom and focuses solely on the peptide backbone. Each amino acid node extends along one of four vectors corresponding to tetrahedral geometry, with unit bond length and a bond angle of approximately $109.4^{\circ}$, consistent with fundamental stereochemical constraints [57]. By removing side-chain variability, the conformational energy surface mapped into the Hamiltonian becomes smoother, reducing high-frequency fluctuations and yielding a formulation that is more robust to the noise characteristics of current quantum devices. The connection vectors between amino acid nodes, inter-residue interactions, and structural constraints are mapped to quantum operators, which together define the system Hamiltonian. The dimensionality of these operators determines the number of qubits required for execution: $n$ qubits span a Hilbert space of dimension $2^{n}$, and thus problem complexity increases with sequence length and amino acid diversity. Table 1 summarizes the number of qubits and circuit depths used in our experiments. For illustration, encoding a residue with four possible extension vectors can be achieved using four qubits, where measurement outcomes such as 0001, 0010, 0100, or 1000 activate one of the four tetrahedral directions [75]. At initialization, all feasible connection paths for a sequence are encoded in superposition, each equally probable. Variational optimization under the Hamiltonian gradually shifts the probability distribution toward low-energy conformations, where invalid configurations are suppressed by energy penalty terms. Upon measurement, the system collapses to a binary string representing the most probable backbone arrangement, corresponding to the conformation that minimizes the protein’s total energy. Through this granularity-controlled encoding, the Hamiltonian retains the essential geometry of peptide chains while discarding details that would otherwise overwhelm current hardware. This design produces a more tractable optimization landscape and enables stable execution of docking-relevant fragments on real quantum processors.
Hamiltonian
Turn encoding and indicator projectors.
For a peptide backbone of length $L$, we allocate 2(L1)2(L-1) qubits, two per turn, to encode the discrete tetrahedral directions. Let the two qubits for turn $t\in\{0,\dots,L-2\}$ be indexed as $(t,0)$ and $(t,1)$. Define single-qubit projectors
$P^{(1)}_{t,b}=\frac{I+Z_{t,b}}{2},\qquad P^{(0)}_{t,b}=\frac{I-Z_{t,b}}{2}\quad(b\in\{0,1\}),$
so that a two-bit direction $d\in\{0,1,2,3\}$ with binary digits $d=(d_{0},d_{1})$ is selected by the two-qubit indicator projector
$\Pi^{(d)}_{t}\;=\;P^{(d_{0})}_{t,0}\,P^{(d_{1})}_{t,1}\;=\;\frac{1}{4}\Big(I+s^{(d)}_{t,0}Z_{t,0}\Big)\Big(I+s^{(d)}_{t,1}Z_{t,1}\Big),$
where $s^{(d)}_{t,b}=+1$ if $d_{b}=1$ and $s^{(d)}_{t,b}=-1$ if $d_{b}=0$. By construction, $\sum_{d=0}^{3}\Pi^{(d)}_{t}=I$ and $\Pi^{(d)}_{t}\Pi^{(e)}_{t}=\delta_{de}\Pi^{(d)}_{t}$. In code, these operators are implemented with SparsePauliOp by composing $\frac{1}{2}(I\pm Z)$ factors for the appropriate qubit indices (see _build_vector_qubit and _build_turn_indicator_fun).
Total Hamiltonian.
The overall Hamiltonian combines all stereochemical, geometric, steric, and interaction constraints:
$H_{\mathrm{t}}\;=\;\lambda_{c}H_{\mathrm{c}}+\lambda_{g}H_{\mathrm{g}}+\lambda_{d}H_{\mathrm{d}}+\lambda_{i}H_{\mathrm{i}},\qquad\lambda_{\bullet}>0.$
Each term reduces to a sparse real linear combination of Pauli strings
$H_{\mathrm{t}}=\sum_{\ell}\gamma_{\ell}\,P_{\ell},\qquad P_{\ell}\in\{I,Z\}^{\otimes n},\;\gamma_{\ell}\in\mathbb{R},$
i.e., an Ising-type Hamiltonian with $Z$-only support, stored compactly in SparsePauliOp. This guarantees that all expectation values can be estimated directly in the computational basis.
(1) Geometric one-hot and turn-consistency constraints.
We enforce that each turn takes exactly one of the four admissible directions and that successive turns respect lattice geometry (bond length and $\sim 109.4^{\circ}$ angle) via
$H_{\mathrm{g}}\;=\;\mu_{\mathrm{oh}}\sum_{t}\Big(1-\sum_{d=0}^{3}\Pi^{(d)}_{t}\Big)^{\!2}\;+\;\mu_{\mathrm{ang}}\sum_{t=1}^{L-2}\sum_{(d,e)\notin\mathcal{A}}\Pi^{(d)}_{t-1}\,\Pi^{(e)}_{t}$,
where $\mathcal{A}$ is the set of allowed direction pairs for adjacent turns under tetrahedral geometry. Expanding the one-hot penalty gives only $I,Z_{t,0},Z_{t,1},Z_{t,0}Z_{t,1}$ terms. The pairwise angle penalties yield at most 4-body $Z$ products.
(2) Chirality and backtracking penalties.
To ensure stereochemical correctness and avoid immediate reversals, we penalize forbidden turn pairs and triples:
$H_{\mathrm{c}}\;=\;\kappa_{\mathrm{bt}}\sum_{t=1}^{L-2}\sum_{(d,e)\in\mathcal{F}_{\mathrm{bt}}}\Pi^{(d)}_{t-1}\,\Pi^{(e)}_{t}\;+\;\kappa_{\chi}\sum_{t=1}^{L-3}\sum_{(d,e,f)\in\mathcal{F}_{\chi}}\Pi^{(d)}_{t-1}\,\Pi^{(e)}_{t}\,\Pi^{(f)}_{t+1}$.
Each triple product expands into up to 6-body $Z$ strings, but remains sparse.
(3) Distance/steric exclusion.
Self-avoidance and steric clashes are encoded via a collision mask $C^{(d,e)}_{t,u}$:
$H_{\mathrm{d}}\;=\;\eta\sum_{0\leq t<u\leq L-2}\;\sum_{d,e=0}^{3}C^{(d,e)}_{t,u}\;\Pi^{(d)}_{t}\,\Pi^{(e)}_{u}$.
This produces sums of 4-body $Z$ strings that penalize invalid spatial contacts.
(4) Pairwise interaction energies.
Residue-residue stabilization is incorporated by a contact mask $M^{(d,e)}_{t,u}$ and interaction matrix $J_{ab}$:
$H_{\mathrm{i}}\;=\;\sum_{0\leq t<u\leq L-2}\;\sum_{d,e=0}^{3}J_{a_{t}a_{u}}\;M^{(d,e)}_{t,u}\;\Pi^{(d)}_{t}\,\Pi^{(e)}_{u}$.
Representative Pauli-$Z$ expansion.
Writing $\Pi^{(d)}_{t}=\frac{1}{4}\sum_{S\subseteq\{(t,0),(t,1)\}}\alpha^{(d)}_{S}\,Z_{S}$ with $Z_{\emptyset}=I$, $Z_{\{(t,b)\}}=Z_{t,b}$ and $Z_{\{(t,0),(t,1)\}}=Z_{t,0}Z_{t,1}$, any quadratic form $\Pi^{(d)}_{t}\Pi^{(e)}_{u}$ expands as
$\Pi^{(d)}_{t}\Pi^{(e)}_{u}=\frac{1}{16}\!\!\sum_{S,T}\alpha^{(d)}_{S}\,\alpha^{(e)}_{T}\;Z_{S}\,Z_{T},$
containing only $I$ and $Z$ products. Thus the full Hamiltonian is an Ising Hamiltonian with sparse Pauli-$Z$ support, directly implementable on utility-level hardware.
Ansatz and optimization.
[…]
@@
In realistic scenarios, the qubit requirement grows significantly with the amino acid sequence length. For very short sequences, fewer than 10 qubits may suffice. However, in our experiments, when dealing with 12 amino acids, at least 77 qubits are needed to encode all relevant operators. Moreover, to satisfy certain coupling conditions and allow for additional SWAP operations that reduce overall circuit depth, we reserve about 5 extra qubits (totaling 82) as a buffer. This strategy effectively lowers the entangling-gate overhead (see, e.g., [41, 40] for related approaches). In many recent studies, the EfficientSU2 ansatz has emerged as a widely used and robust parameterized circuit, recognized for its relatively broad applicability and moderate resource requirements [26, 15]. To parameterize our circuits, we employ the EfficientSU2 ansatz from Qiskit. This ansatz typically consists of repeated blocks of single-qubit rotation gates $R_{y}$ and $R_{z}$, interspersed with CX gates arranged in a circular or linear topology to introduce entanglement. By adjusting the number of repetition layers (reps), we can flexibly control the circuit’s expressive power. In a single repetition layer, each qubit undergoes one $R_{y}R_{z}$ rotation, followed by an entangling step that couples adjacent qubits. The theoretical circuit depth of Qiskit’s EfficientSU2 parameterized quantum circuit grows linearly with the number of qubits $N$. For our case with $N=82$ qubits: Total theoretical depth is $333$. The actual compiled depth may be different through gate optimization and hardware topology adaptation. An important advantage of the quantum method is its extremely fast computational speed. Due to the superposition state of the quantum system, during each optimization, the system does not focus on a specific structure for validation like traditional structure-based prediction methods. Instead, the superposition state allows all possible conformations of the current protein structure to exist simultaneously at the physical level, collapsing into a definite ground state as the optimal result through parameter optimization. In our tests, for a sequence with a length of 5 amino acids, the total recorded execution time after 150 iterations was 4079.55 seconds. According to logs from IBM’s superconducting quantum computer, the average time per optimization iteration and state measurement, from task submission to result acquisition, was approximately 10 seconds, with about 73.53% of the time spent on the quantum end. In our setup, the number of shots for optimization tasks was 200, and the number of shots for measurement was 100,000, indicating that the execution efficiency of the quantum computing method is extremely high. The quantum processor itself completes a single task almost instantaneously, with most of the time spent on state preparation and the machine’s self-calibration. Additionally, part of the time was limited by the performance of the classical computing backend. Under computational models of equivalent accuracy, the quantum method is several orders of magnitude faster than all current methods based on classical computing.